# A Comprehensive Survey of Scientific Large Language Models and Their Applications in Scientific Discovery

### v1

@@ 5.2 Language + Graph
In some scenarios, molecular graphs appear simultaneously with text information, thus existing works have explored how to encode both effectively. The first type of such models adopt a GNN as the graph encoder and an LLM as the text encoder. The two modalities are connected through contrastive learning Liu et al. (2023d) (Type 3.c). For example, Text2Mol Edwards et al. (2021) uses GCN Kipf and Welling (2016) and SciBERT to encode a molecule and its corresponding natural language description, respectively, for text-to-molecule retrieval. The second type of models utilize an LLM to encode text and graphs simultaneously Zeng et al. (2022). Graphs can be either linearized to SMILES strings Edwards et al. (2022) (Type 2.c) or projected onto virtual tokens with graph encoders Zhao et al. (2023a); Liu et al. (2023f) (Type 2.d). For instance, 3D-MoLM Li et al. (2024b) uses a 3D molecular encoder to represent molecules as tokens, and feed them together with instructions into LLaMA-2 for molecule-to-text retrieval and molecule captioning.
@@ 5.3 Language + Vision
Complementing text and graph modalities, molecular images form the vision modality in chemistry. Existing works adopt a similar philosophy to BLIP-2 Li et al. (2023b), which represents each image as tokens and feed them into an LLM (Type 2.d). For example, GIT-Mol Liu et al. (2024) projects all modalities, including graphs and images, into the latent text space and conducts encoding and decoding with T5 Raffel et al. (2020).
@@ 5.4 Molecule
Different from subsection 5.2, this subsection introduces models dealing with molecules without associated text information. That being said, comparable approaches inspired by LLMs are utilized to develop molecular language models Flam-Shepherd et al. (2022). To be specific, most studies adopt SMILES or SELFIES Krenn et al. (2020) strings as the sequential representation of molecules. Similar to the trend in the “Language” modality, pioneering molecular LLMs focus on representation learning with bidirectional Transformer encoders (Type 1.c, e.g., SMILES-BERT Wang et al. (2019) and MoLFormer Ross et al. (2022)). For instance, ChemBERTa Chithrananda et al. (2020) adopts the architecture and pre-training strategy similar with those of RoBERTa Liu et al. (2019). These models exhibit extraordinary abilities in molecular understanding tasks such as molecular property prediction (e.g., toxicity classification Wu et al. (2018) and atomization energy regression Ramakrishnan et al. (2014)) as well as virtual screening Riniker and Landrum (2013). Later works explore representing molecules in an autoregressive fashion (Type 2.c, e.g., BARTSmiles Chilingaryan et al. (2022) and ChemGPT Frey et al. (2023)). For instance, T5Chem Lu and Zhang (2022) adopts the T5 backbone and a sequence-to-sequence pre-training objective. These models are evaluated in generative tasks that include molecule generation Gaulton et al. (2017), reaction prediction, and retrosynthesis. Besides linearizing molecules, there are studies modifying the Transformer architecture to admit molecular graphs, such as MAT Maziarka et al. (2020) and R-MAT Maziarka et al. (2024).
@@ 5.5 Applications in Scientific Discovery
Previous studies have shown that LLMs facilitate autonomous chemical research. For example, Bran et al. (2024) present a chemistry LLM agent, ChemCrow, that can integrate expert-designed tools for organic synthesis, drug discovery, and materials design; Boiko et al. (2023) develop an LLM-empowered intelligence system, Coscientist, that can design, plan, and perform chemical research. LLMs also help with drug and catalyst design. For instance, ChatDrug Liu et al. (2023e) explores drug editing using LLMs with a prompt module, a domain feedback module, and a conversation module; DrugAssist Ye et al. (2023a) is proposed as an LLM-based interactive model for molecule optimization through human-machine dialogue; Sprueill et al. (2023, 2024) use LLMs as agents to search for effective catalysts through Monte Carlo Tree Search and the feedback from an atomistic neural network model.
@@ 6.1 Language
Besides research articles (e.g., titles/abstracts from PubMed Lu (2011) and full text from PMC Beck and Sequeira (2003)), pre-training corpora for biomedical LLMs include electronic health records (e.g., MIMIC-III Johnson et al. (2016), MIMIC-IV Johnson et al. (2023)), knowledge bases (e.g., UMLS Bodenreider (2004)), and health-related social media posts (e.g., COVID-19 tweets Müller et al. (2023)). Recent studies further collect supervised fine-tuning and preference optimization datasets from medical exam questions, knowledge graphs, and doctor-patient dialogues. Examples include ChiMed Ye et al. (2023b), MedInstruct-52k Zhang et al. (2023e), and BiMed1.3M Acikgoz et al. (2024), many of which have non-English components (e.g., Chinese and Arabic).
The watershed moment of biomedical LLM evolution is still the emergence of billion-parameter architectures and instruction tuning. Before that, a wide variety of moderate-sized backbones are explored, including both encoder-based (Type 1.a, e.g., BioBERT Lee et al. (2020), Bio-ELECTRA Ozyurt (2020), BioRoBERTa Lewis et al. (2020b), BioALBERT Naseem et al. (2022), and Clinical-Longformer Li et al. (2022a)) and (encoder-) decoder-based ones (Type 2.a, e.g., SciFive Phan et al. (2021), BioBART Yuan et al. (2022a), and BioGPT Luo et al. (2022)). Evaluation tasks for these models range from biomedical NER, RE, sentence similarity estimation, document classification, and QA (i.e., the BLURB bechmark Gu et al. (2021)) to natural language inference (NLI) Romanov and Shivade (2018) and entity linking Doğan et al. (2014). After the watershed, the trend becomes instruction-tuning billion-parameter LLMs (Type 2.a, e.g., Med-PaLM Singhal et al. (2023a), MedAlpaca Han et al. (2023), and BioMistral Labrak et al. (2024)). Accordingly, evaluation tasks become single-round QA Jin et al. (2021); Pal et al. (2022) and multi-round dialogue Wang et al. (2023h). Meanwhile, there are studies proposing a Bi-Encoder architecture (Type 3.a, e.g., Jin et al. (2023c) and Xu et al. (2024)) that specifically targets biomedical retrieval tasks, the benchmarks of which are NFCorpus Boteva et al. (2016), TREC-COVID Voorhees et al. (2021), etc.
@@ 6.2 Language + Graph
Biomedical ontologies capture rich types of relations between entities. Analogously, citation links characterize connections between biomedical papers. Intuitively, jointly leveraging text and such graph information paves the way for multi-hop reasoning in QA. For instance, Yasunaga et al. (2022a) propose to use an LLM and a GNN to encode text and ontology signals, respectively, and deeply fuse them (Type 3.c); Yasunaga et al. (2022b) concatenate text segments from two linked papers together and feed the sequence into an LLM for pre-training, which is essentially appending a metadata neighbor (i.e., reference) as context for MLM (Type 1.b). Both approaches demonstrate significant improvement on QA tasks that require complex reasoning.
@@ 6.3 Language + Vision
Most biomedical vision-language models exploit the CLIP architecture Radford et al. (2021), where a text encoder and an image encoder is jointly trained to map the paired text and image closer via contrastive learning (Type 3.d). The choice of the text encoder evolves from BERT Zhang et al. (2022) and GPT-2 Huang et al. (2023b) to LLaMA Wu et al. (2023) and LLaMA-2 Liu et al. (2023b), while the image encoder evolves from ResNet Huang et al. (2021) to ViT Zhang et al. (2023c) and Swin Transformer Thawkar et al. (2023). MLM, masked image modeling, and text-text/image-image contrastive learning (i.e., by creating augmented views within the language/vision modality) are sometimes adopted as auxiliary pre-training tasks. Besides CLIP, other general-domain vision-language architectures, such as LLaVA Li et al. (2023a), PaLM-E Tu et al. (2024), and Gemini Saab et al. (2024), have been explored. For instance, LLaVA-Med (Type 2.d) encodes images onto several visual tokens and prepends them to text tokens as the LLM input. Evaluation tasks of these models encompass image classification, segmentation, object detection, vision QA, text-to-image/image-to-text retrieval, and report generation, the benchmarks of which include CheXpert Irvin et al. (2019), PadChest Bustos et al. (2020), SLAKE Liu et al. (2021a), etc.
@@ 6.5 Applications in Scientific Discovery
Similarly to chemistry, LLMs can automate experiments in biology and medicine research. For example, CRISPR-GPT Huang et al. (2024) augments an LLM agent with domain knowledge to enhance the design process of CRISPR-based gene-editing experiments. Moreover, LLMs can encode biological sequences to capture structural properties, guide protein design, and evaluate the evolutionary fitness of viral variants. For instance, ESM-1b Rives et al. (2021) and ESM-2 Lin et al. (2023b) enable accurate structure prediction of proteins without expensive and time-consuming experiments; Ferruz and Höcker (2022) fine-tune LLMs on protein families, which can generate highly divergent but still potentially functional novel sequences; Hie et al. (2021) develop LLMs that can predict viral escape mutations.
@@ 7.1 Language
Geoscience research papers, climate-related news articles, Wikipedia pages, corporate sustainability reports, knowledge bases (e.g., GAKG Deng et al. (2021)), and point-of-interest (POI) data (e.g., OpenStreetMap Haklay and Weber (2008)) constitute the pre-training corpora of geoscience LLMs.
Preliminary research on geoscience LLMs focuses on pre-training bidirectional LLMs with the Transformer encoder backbone (Type 1.a, e.g., ClimateBERT Webersinke et al. (2021), SpaBERT Li et al. (2022b), and MGeo Ding et al. (2023)). For instance, SpaBERT and MGeo perform MLM on a sequence of geolocations for geographic entity linking and query-POI matching, respectively. More recently, related studies concentrate on scaling up decoding-style autoregressive LLMs in geoscience (Type 2.a, e.g., K2 Deng et al. (2024), OceanGPT Bi et al. (2023b), and GeoGalactica Lin et al. (2024b)). For instance, K2 and OceanGPT adapt LLaMA to geoscience and ocean science, respectively, via supervised fine-tuning with domain-specific instructions curated by human experts and/or augmented by general-domain LLMs. Evaluations of such models are conducted on geoscience benchmarks, such as GeoBench Deng et al. (2024) and OceanBench Bi et al. (2023b), which contain a broad range of tasks including QA, classification, knowledge probing, reasoning, summarization, and generation.
@@ 7.2 Language + Graph
Some geoscience applications involve graph signals, such as heterogeneous POI networks and knowledge graphs. To handle such signals and text jointly, ERNIE-GeoL Huang et al. (2022) introduces a transformer-based aggregation layer to deeply fuse text and POI information within the BERT-based architecture; PK-Chat Deng et al. (2023) combines an LLM with a pointer generation network on a knowledge graph to build a knowledge-driven dialogue system.
@@ 7.3 Language + Vision
Aerial views, together with location descriptions, profile urban regions. To deal with language and vision modalities jointly, UrbanCLIP Yan et al. (2024) considers the CLIP architecture (Type 3.d), which is also widely adopted by biomedical vision-language models as mentioned in subsection 6.3, to perform text-image contrastive learning for urban indicator prediction.
@@ 7.5 Applications in Scientific Discovery
In geography, Wang et al. (2023b) and Zhou et al. (2024) highlight the potential of LLMs in urban planning from the sustainability, living, economic, disaster, and environmental perspectives. In geology, besides climate and weather forecasting, foundation models have been applied to simultaneous earthquake detection and phase picking Mousavi et al. (2020). In environmental science, ChatClimate Vaghefi et al. (2023) enhances GPT-4 by providing access to external, scientifically accurate knowledge on climate change to build a climate science conversational AI.
@@ 8 Challenges and Future Directions
Diving into Fine-Grained Themes. Most existing scientific LLMs target a coarse-grained field (e.g., chemistry), while some tasks rely on highly specialized knowledge of a fine-grained theme (e.g., Suzuki coupling). When LLMs are pre-trained on more general corpora, frequently appeared signals may dominate the model parameter space, and domain-specific tail knowledge may swipe out. We believe automatically curating in-depth, theme-focused knowledge graphs Hope et al. (2021) and using them to guide the generation process will be a promising direction to tackle this issue.
Generalizing to Out-of-Distribution Scientific Data. In the scientific domain, it is common that the testing distribution shifts from the training distribution Zhang et al. (2023f): novel scientific concepts keep emerging in newly published papers; unseen molecules with different scaffolds and unseen proteins with different numbers of peptide chains may appear during testing. Handling such out-of-distribution data remains a challenge for pre-trained scientific LLMs. To our knowledge, invariant learning Arjovsky et al. (2019) can serve as the theoretical foundation for out-of-distribution analyses, and how to integrate it into LLM pre-training is worth exploring.
Facilitating Trustworthy Predictions. LLMs can generate plausible-sounding but factually incorrect output, commonly known as hallucination Ji et al. (2023), which is particularly dangerous in high-stake scientific domains such as chemistry and biomedicine. To mitigate this issue, retrieval-augmented generation (RAG) provides LLMs with relevant, up-to-date, and trustworthy information. However, previous RAG studies in the scientific domain mainly focus on retrieving text Xiong et al. (2024) and knowledge Jin et al. (2024), while scientific data are heterogeneous and multi-modal. We envision that cross-modal RAG (e.g., guiding text generation with relevant chemicals and proteins) will present additional opportunities to further enhance the trustworthiness of scientific LLMs.
@@
This survey mainly covers LLMs in mathematics and natural sciences. We are aware that LLMs can also significantly impact social sciences by achieving remarkable performance in representative tasks Ziems et al. (2024) and serving as agents for social simulation experiments Horton (2023), but we leave the survey of these efforts as future work due to space limitations. In addition, this paper focuses on LLMs pre-trained on scientific data or augmented with domain-specific knowledge to benefit scientific discovery. There are studies Wang et al. (2023g); Guo et al. (2023) proposing new benchmark datasets of scientific problems but evaluating the performance of general-purpose LLMs only, and we do not include these works in our survey. Furthermore, some LLMs may belong to more than one field or modality category given our classification criteria in the paper. For instance, BioMedGPT Luo et al. (2023c) is pre-trained on biology and chemistry data jointly; GIT-Mol Liu et al. (2024) considers the language, graph, and vision modalities simultaneously. For succinctness, we introduce each of them in only one subsection.

### v2

@@ 5.2 Language + Graph
In some scenarios, molecular graphs appear simultaneously with text information, thus existing works have explored how to encode both effectively. The first type of such models adopts a GNN as the graph encoder and an LLM as the text encoder. The two modalities are connected through contrastive learning Liu et al. (2023d) (Type 3.c). For example, Text2Mol Edwards et al. (2021) uses GCN Kipf and Welling (2016) and SciBERT to encode a molecule and its corresponding natural language description, respectively, for text-to-molecule retrieval. The second type of models utilizes an LLM to encode text and graphs simultaneously Zeng et al. (2022). Graphs can be either linearized to SMILES strings Edwards et al. (2022) (Type 2.c) or projected onto virtual tokens with graph encoders Zhao et al. (2023a); Liu et al. (2023e) (Type 2.d). For instance, 3D-MoLM Li et al. (2024c) uses a 3D molecular encoder to represent molecules as tokens, and feeds them together with instructions into LLaMA-2 for molecule-to-text retrieval and molecule captioning.
@@ 5.3 Language + Vision
Complementing text and graph modalities, molecular images form the vision modality in chemistry. Existing works adopt a similar philosophy to BLIP-2 Li et al. (2023b), which represents each image as tokens and feeds them into an LLM (Type 2.d). For example, GIT-Mol Liu et al. (2024a) projects all modalities, including graphs and images, into the latent text space and conducts encoding and decoding using T5 Raffel et al. (2020).
@@ 5.4 Molecule
Different from subsection 5.2, this subsection introduces models dealing with molecules without associated text information. That being said, comparable approaches inspired by LLMs are utilized to develop molecular language models Flam-Shepherd et al. (2022). To be specific, most studies adopt SMILES or SELFIES Krenn et al. (2020) strings as the sequential representation of molecules. Similar to the trend in the “Language” modality, pioneering molecular LLMs focus on representation learning with bidirectional Transformer encoders (Type 1.c, e.g., SMILES-BERT Wang et al. (2019) and MoLFormer Ross et al. (2022)). For instance, ChemBERTa Chithrananda et al. (2020) adopts the architecture and pre-training strategy similar to those of RoBERTa Liu et al. (2019). These models exhibit extraordinary abilities in molecular understanding tasks such as molecular property prediction (e.g., toxicity classification Wu et al. (2018) and atomization energy regression Ramakrishnan et al. (2014)) as well as virtual screening Riniker and Landrum (2013). Later works explore representing molecules in an autoregressive fashion (Type 2.c, e.g., BARTSmiles Chilingaryan et al. (2024) and ChemGPT Frey et al. (2023)). For instance, T5Chem Lu and Zhang (2022) adopts the T5 backbone and a sequence-to-sequence pre-training objective. These models are evaluated in generative tasks that include molecule generation Gaulton et al. (2017), reaction prediction, and retrosynthesis. Besides linearizing molecules, there are studies modifying the Transformer architecture to admit molecular graphs, such as MAT Maziarka et al. (2020) and R-MAT Maziarka et al. (2024).
@@ 5.5 Applications in Scientific Discovery
Previous studies have shown that LLMs facilitate autonomous chemical research. For example, Bran et al. (2024) present a chemistry LLM agent, ChemCrow, that can integrate expert-designed tools for organic synthesis, drug discovery, and materials design; Zheng et al. (2023a) demonstrate that LLMs can perform knowledge synthesis from the scientific literature, knowledge inference from data, and interpretable explanation generation in chemistry; Boiko et al. (2023) develop an LLM-empowered intelligence system, Coscientist, that can design, plan, and perform chemical research. Moreover, LLMs accomplish complex tasks in chemistry, such as drug and catalyst design and molecular discovery, purely from instructions White (2023). For instance, Ramos et al. (2023) study catalyst and molecule design with in-context learning, removing the requirement for traditional training or simulation processes; ChatDrug Liu et al. (2024b) explores drug editing using LLMs with a prompt module, a domain feedback module, and a conversation module; Jablonka et al. (2024) find that fine-tuned LLMs perform comparably to, or even better than, conventional techniques for many chemistry applications, spanning from the properties of molecules and materials to the yield of chemical reactions; DrugAssist Ye et al. (2023a) serves as an LLM-based interactive model for molecule optimization through human-machine dialogue; Sprueill et al. (2023, 2024) use LLMs as agents to search for effective catalysts through Monte Carlo Tree Search and the feedback from an atomistic neural network model; Wang et al. (2024b) re-engineer crossover and mutation operations for molecular discovery using LLMs trained on extensive chemical datasets. Meanwhile, benchmarking studies by Mirza et al. (2024) demonstrate that although LLMs achieve superhuman proficiency in many chemical tasks, further research is critical to enhancing their safety and utility in chemical sciences.
@@ 6.1 Language
Besides research articles (e.g., titles/abstracts from PubMed Lu (2011) and full text from PMC Beck and Sequeira (2003)), pre-training corpora for biomedical LLMs include electronic health records (e.g., MIMIC-III Johnson et al. (2016), MIMIC-IV Johnson et al. (2023)), knowledge bases (e.g., UMLS Bodenreider (2004)), and health-related social media posts (e.g., COVID-19 tweets Müller et al. (2023)). Recent studies further collect supervised fine-tuning and preference optimization datasets from medical exam questions, knowledge graphs, and doctor-patient dialogues. Examples include ChiMed Ye et al. (2023b), MedInstruct-52k Zhang et al. (2023d), and BiMed1.3M Acikgoz et al. (2024), many of which have non-English components (e.g., Chinese and Arabic).
The watershed moment in the evolution biomedical LLMs is still the emergence of billion-parameter architectures and instruction tuning. Before that, a wide variety of moderate-sized backbones are explored, including both encoder-based (Type 1.a, e.g., BioBERT Lee et al. (2020), Bio-ELECTRA Ozyurt (2020), BioRoBERTa Lewis et al. (2020b), BioALBERT Naseem et al. (2022), and Clinical-Longformer Li et al. (2022a)) and (encoder-) decoder-based ones (Type 2.a, e.g., SciFive Phan et al. (2021), BioBART Yuan et al. (2022a), and BioGPT Luo et al. (2022)). Evaluation tasks for these models range from biomedical NER, RE, sentence similarity estimation, document classification, and QA (i.e., the BLURB benchmark Gu et al. (2021)) to natural language inference (NLI) Romanov and Shivade (2018) and entity linking Doğan et al. (2014). After the watershed, the trend becomes instruction-tuning billion-parameter LLMs (Type 2.a, e.g., Med-PaLM Singhal et al. (2023a), MedAlpaca Han et al. (2023), and BioMistral Labrak et al. (2024)). Accordingly, evaluation tasks now include single-round QA Jin et al. (2021); Pal et al. (2022) and multi-round dialogue Wang et al. (2023f). Meanwhile, there are studies proposing a Bi-Encoder architecture (Type 3.a, e.g., Jin et al. (2023c) and Xu et al. (2024)) that specifically targets biomedical retrieval tasks, the benchmarks of which are NFCorpus Boteva et al. (2016), TREC-COVID Voorhees et al. (2021), etc.
@@ 6.2 Language + Graph
Biomedical ontologies capture rich types of relations between entities. Analogously, citation links characterize connections between biomedical papers. Intuitively, jointly leveraging text and such graph information paves the way for multi-hop reasoning in QA. For instance, Yasunaga et al. (2022a) propose to use an LLM and a GNN to encode text and ontology signals, respectively, and deeply fuse them (Type 3.c); Yasunaga et al. (2022b) concatenate text segments from two linked papers together and feed the sequence into an LLM for pre-training, which is essentially appending a metadata neighbor (i.e., reference) as context for MLM (Type 1.b). Both approaches demonstrate significant improvement in QA tasks that require complex reasoning.
@@ 6.3 Language + Vision
Most biomedical vision-language models exploit the CLIP architecture Radford et al. (2021), where a text encoder and an image encoder are jointly trained to map the paired text and image closer via contrastive learning (Type 3.d). The choice of the text encoder evolves from BERT Zhang et al. (2022) and GPT-2 Huang et al. (2023b) to LLaMA Wu et al. (2023) and LLaMA-2 Liu et al. (2023b), while the image encoder evolves from ResNet Huang et al. (2021) to ViT Zhang et al. (2023c) and Swin Transformer Thawkar et al. (2024). MLM, masked image modeling, and text-text/image-image contrastive learning (i.e., by creating augmented views within the language/vision modality) are sometimes adopted as auxiliary pre-training tasks. Besides CLIP, other general-domain vision-language architectures, such as LLaVA Li et al. (2023a), PaLM-E Tu et al. (2024), and Gemini Saab et al. (2024), have been explored. For instance, LLaVA-Med (Type 2.d) encodes images onto several visual tokens and prepends them to text tokens as the LLM input. Evaluation tasks of these models encompass image classification, segmentation, object detection, vision QA, text-to-image/image-to-text retrieval, and report generation, the benchmarks of which include CheXpert Irvin et al. (2019), PadChest Bustos et al. (2020), SLAKE Liu et al. (2021a), etc.
@@ 6.5 Applications in Scientific Discovery
Similarly to chemistry, LLMs can automate experiments in biological and medical research. For example, CRISPR-GPT Huang et al. (2024) augments an LLM agent with domain knowledge to enhance the design process of CRISPR-based gene-editing experiments; TrialMind Wang et al. (2024f) utilizes LLMs to extract results and synthesize clinical evidence from the literature for medical discovery. Moreover, LLMs can encode biological sequences to capture structural properties, guide protein design, and evaluate the evolutionary fitness of viral variants. For instance, ESM-1b Rives et al. (2021) and ESM-2 Lin et al. (2023b) enable accurate structure prediction of proteins without expensive and time-consuming experiments; Ferruz and Höcker (2022) fine-tune LLMs on protein families, which can generate highly divergent but still potentially functional novel sequences; Hie et al. (2021) develop LLMs that can predict viral escape mutations.
@@ 7.1 Language
Geoscience research papers, climate-related news articles, Wikipedia pages, corporate sustainability reports, knowledge bases (e.g., GAKG Deng et al. (2021)), and point-of-interest (POI) data (e.g., OpenStreetMap Haklay and Weber (2008)) constitute the pre-training corpora for geoscience LLMs.
Preliminary research on geoscience LLMs focuses on pre-training bidirectional LLMs with the Transformer encoder backbone (Type 1.a, e.g., ClimateBERT Webersinke et al. (2021), SpaBERT Li et al. (2022b), and MGeo Ding et al. (2023)). For instance, SpaBERT and MGeo perform MLM on a sequence of geolocations for geographic entity linking and query-POI matching, respectively. More recently, related studies concentrate on scaling up decoding-style autoregressive LLMs in geoscience (Type 2.a, e.g., K2 Deng et al. (2024), OceanGPT Bi et al. (2023b), and GeoGalactica Lin et al. (2024c)). For instance, K2 and OceanGPT adapt LLaMA to geoscience and ocean science, respectively, via supervised fine-tuning with domain-specific instructions curated by human experts and/or augmented by general-domain LLMs. Evaluations of such models are conducted on geoscience benchmarks, such as GeoBench Deng et al. (2024) and OceanBench Bi et al. (2023b), which encompass a broad range of tasks including QA, classification, knowledge probing, reasoning, summarization, and generation.
@@ 7.2 Language + Graph
Some geoscience applications involve graph signals, such as heterogeneous POI networks and knowledge graphs. To handle such signals and text jointly, ERNIE-GeoL Huang et al. (2022) introduces a Transformer-based aggregation layer to deeply fuse text and POI information within a BERT-based architecture; PK-Chat Deng et al. (2023) combines an LLM with a pointer generation network on a knowledge graph to build a knowledge-driven dialogue system.
@@ 7.3 Language + Vision
Aerial views, together with location descriptions, profile urban regions. To address language and vision modalities jointly, UrbanCLIP Yan et al. (2024) considers the CLIP architecture (Type 3.d), which is also widely adopted by biomedical vision-language models as mentioned in subsection 6.3, to perform text-image contrastive learning for urban indicator prediction.
@@ 7.5 Applications in Scientific Discovery
In geography, Wang et al. (2023b) and Zhou et al. (2024b) highlight the potential of LLMs in urban planning from sustainability, living, economic, disaster, and environmental perspectives. In geology, besides climate and weather forecasting, foundation models have been applied to simultaneous earthquake detection and phase picking Mousavi et al. (2020). In environmental science, ChatClimate Vaghefi et al. (2023) enhances GPT-4 by providing access to external, scientifically accurate knowledge on climate change to build a climate science conversational AI.
@@ 8 Challenges and Future Directions
Diving into Fine-Grained Themes. Most existing scientific LLMs target a coarse-grained field (e.g., chemistry), while some tasks rely on highly specialized knowledge of a fine-grained theme (e.g., Suzuki coupling). When LLMs are pre-trained on more general corpora, frequently appeared signals may dominate the model parameter space, and domain-specific tail knowledge may be wiped out. We believe that automatically curating in-depth, theme-focused knowledge graphs Hope et al. (2021) to guide the generation process will be a promising direction to tackle this issue.
Generalizing to Out-of-Distribution Scientific Data. In the scientific domain, it is common that the testing distribution shifts from the training distribution Zhang et al. (2023e): novel scientific concepts keep emerging in newly published papers; unseen molecules with different scaffolds and unseen proteins with different numbers of peptide chains may appear during testing. Handling such out-of-distribution data remains a challenge for pre-trained scientific LLMs. To our knowledge, invariant learning Arjovsky et al. (2019) can serve as the theoretical foundation for out-of-distribution analyses, and how to integrate it into LLM pre-training is worth exploring.
Facilitating Trustworthy Predictions. LLMs can generate plausible-sounding but factually incorrect output, commonly known as hallucination Ji et al. (2023), which is particularly dangerous in high-stakes scientific domains such as chemistry and biomedicine. To mitigate this issue, retrieval-augmented generation (RAG) provides LLMs with relevant, up-to-date, and trustworthy information. However, previous RAG studies in the scientific domain mainly focus on retrieving text Xiong et al. (2024) and knowledge Jin et al. (2024), while scientific data are heterogeneous and multi-modal. We envision that cross-modal RAG (e.g., guiding text generation with relevant chemicals and proteins) will present additional opportunities to further enhance the trustworthiness of scientific LLMs.
@@
This survey primarily covers LLMs in mathematics and natural sciences. We are aware that LLMs can also significantly impact social sciences by achieving remarkable performance in representative tasks Ziems et al. (2024) and serving as agents for social simulation experiments Horton (2023), but we leave the survey of these efforts as future work due to space limitations. In addition, this paper focuses on LLMs pre-trained on scientific data or augmented with domain-specific knowledge to benefit scientific discovery. There are studies Guo et al. (2023); Wang et al. (2024e); Yue et al. (2024a); Liang et al. (2024d) proposing new benchmark datasets of scientific problems but evaluating the performance of general-purpose LLMs only, and we do not include these works in our survey. Furthermore, some LLMs may belong to more than one field or modality category given our classification criteria in the paper. For instance, BioMedGPT Luo et al. (2023c) is pre-trained on biology and chemistry data jointly; GIT-Mol Liu et al. (2024a) considers the language, graph, and vision modalities simultaneously. For the sake of brevity, we introduce each of them in only one subsection.